# Variable Neighborhood Searching Rerandomization

### v1

@@ 3.2 Statistical Properties of VNSRR
And we focus on three key properties of rerandomization algorithms.
Property 1
(symmetric feasible set) $\mathcal{F}=1-\mathcal{F}$.
Property 2
(uniform initialization) For any $w\in\mathcal{F}$,
$P\left(W^{(0)}=w\right)=\frac{1}{|\mathcal{F}|}.$
Property 3
$P\left(W^{(t)}=w^{(t)}\mid W^{(0)}=w^{(0)},\ldots,W^{(t-1)}=w^{(t-1)}\right)=u_{t}(M(w^{(0)}),\ldots,M(w^{(t)})),$
Based on Property 1-3, we prove that the difference-in-means estimator is unbiased.
Theorem 1
If $W$ is generated by a rerandomization algorithm satisfying Property 1, Property 2 and Property 3, then $E[\widehat{\tau}]=\tau$.
Based on Condition 1-2 and Property 1-3, we prove that the difference-in-means estimator has smaller variance than that under complete randomization and construct a lower bound for the variance reduction. Denote $V_{CR}$ as the variance of the difference-in-means estimator under complete randomization, we prove the following theorem.
Theorem 2
Under Condition 1 and Condition 2, if $W$ is generated by a rerandomization algorithm satisfying Property 1, Property 2 and Property 3, then
$\frac{V_{CR}-Var[\widehat{\tau}]}{V_{CR}}\geq\left(1-\frac{a}{p}\right)R^{2},$
where $R^{2}=\frac{\beta^{T}Cov_{CR}[\overline{X}_{t}-\overline{X}_{c}]\beta}{V_{CR}}=\frac{n\beta^{T}S_{XX}\beta}{n_{t}n_{c}V_{CR}}$. Similar to the definition of $V_{CR}$, $Cov_{CR}[\overline{X}_{t}-\overline{X}_{c}]$ represents the covariance matrix of $\overline{X}_{t}-\overline{X}_{c}$ under complete randomization.
One can verify that VNSRR with $\mathcal{F}_{0}$ as the feasible set satisfies our proposed requirements under the following conditions.
Condition 3
@@
Proposition 1
VNSRR satisfies Property 2 and Property 3. Under Condition 3, VNSRR satisfies Property 1.
Therefore, the unbiasedness and variance reduction of the difference-in-means estimator under VNSRR are immediately evident as corollaries of Theorem 1 and Theorem 2.
Corollary 1
Under Condition 3 and VNSRR, $E[\widehat{\tau}]=\tau$.
Corollary 2
Under Condition 1, Condition 2, Condition 3, and VNSRR,
$\frac{V_{CR}-Var[\widehat{\tau}]}{V_{CR}}\geq\left(1-\frac{a}{p}\right)R^{2}.$
One can verify that CR, ARSRR, and PSRR all satisfy Property 1, Property 2, and Property 3 under certain conditions. As a result, their statistical properties are established by this framework as well. All verifications and proofs are included in the Supplement.
4 VNSRR for Structured Randomized Experiments
In this section, we extend VNSRR to randomized experiments with various types of structures. For each VNSRR variant, we derive the unbiasedness and a lower bound for the variance reduction of the difference-in-means estimator. We will reuse several notations across different settings, leveraging the similarities between settings to streamline our notation system. The pseudocode of each variant, along with the theoretical results and their proofs, is provided in the Supplement.
4.1 VNSRR for Sequentially Randomized Experiments
In Section 3, we consider offline rerandomization, which assigns treatment to all units at once. However, as pointed out by Zhou et al. (2018), experimenters may be unable to wait for all units to arrive before performing the experiments. It is common to assign treatment to units group by group in practice. To balance the covariates in such sequential experiments, Zhou et al. (2018) suggested performing sequential rerandomization, which can be computationally challenging. Bertsimas et al. (2019) proposed a novel algorithm for online allocation that leverages robust mixed-integer optimization. However, the authors focused on optimal designs, which are inherently different from the rerandomization framework.
@@ 4.1 VNSRR for Sequentially Randomized Experiments
We now review the basic concepts and notations of sequential rerandomization. We assume that the experiment consists of $K$ stages, and the $n$ units are divided into $K$ sequential groups of sizes $(n_{1},\ldots,n_{K})$. In the $k$-th stage, $n_{tk}$ units in the $k$-th group are assigned to the treatment arm, and the remaining $n_{ck}$ units are assigned to the control arm. We use the subscript $[k]$ to represent the reduced version of a term calculated based on only the first $k$ groups. The covariates of the first $k$ groups are denoted by $X_{[k]}$ and its covariance matrix is denoted by $S_{XX[k]}$. The assignment status of the first $k$ groups is denoted by $W_{[k]}$. And we define $n_{[k]}=\sum_{g=1}^{k}n_{g}$, $n_{t[k]}=\sum_{g=1}^{k}n_{tg}$, $n_{c[k]}=\sum_{g=1}^{k}n_{cg}$, $H_{[k]}=\frac{n_{[k]}}{n_{t[k]}n_{c[k]}}X_{[k]}S_{XX[k]}^{-1}X_{[k]}^{T}$. The Mahalanobis distance corresponding to $W_{[k]}$ is defined as
$M_{[k]}$ $=(W_{[k]}-\frac{n_{t[k]}}{n_{[k]}}1_{n_{[k]}})^{T}H_{[k]}(W_{[k]}-\frac{n_{t[k]}}{n_{[k]}}1_{n_{[k]}})$
Sequential rerandomization proceeds as follows: When the $k$-th group enrolls in the experiment, the experimenter randomly assigns treatment to the units without changing the assignment status of the previously enrolled groups. The experimenter aims to balance the covariates of all units in the first $k$ groups. Therefore, the experimenter rerandomizes the $n_{k}$ units until $M_{[k]}\leq a_{k}$, where $a_{k}$ is a prespecified threshold. Zhou et al. (2018) proposed to use acceptance-rejection sampling in this procedure, so we refer to their method as SeqARSRR. Recently, Zhu and Liu (2022) proposed a sequential pair-switching rerandomization(SeqPSRR) method to reduce the computational cost. SeqPSRR essentially performs PSRR in each stage of the sequential experiment. In Section 5.2, we will show that both SeqARSRR and SeqPSRR can be time-consuming.
We note that sequential rerandomization in the $k$-th stage can be formulated into the following constrained binary quadratic programming problem, whose feasible set we refer to as $\mathcal{F}_{[k]}$.
minimize $\quad M_{[k]}(W_{[k]}),$
subject to $\quad\sum_{i=n_{[k-1]}+1}^{n_{[k]}}W_{[k],i}=n_{tk},$
$\quad W_{[k],i}\in\{0,1\},\;i=n_{[k-1]}+1,\ldots,n_{[k]},$
$\quad W_{[k],i}=W_{[k-1],i},\;i=1,\ldots,n_{[k-1]}.$ (2)
Based on this observation, we extend VNSRR to sequentially randomized experiments and refer to this variant as SeqVNSRR. SeqVNSRR performs VNSRR in each stage of the experiment while fixing the assignments obtained in previous stages. It also adopts the updating trick, the specific form of the local search step, and the hyperparameter combination mentioned in Section 3 to reduce the computational cost.
To analyze the statistical properties of SeqVNSRR, we consider the following variants of Condition 3 and Property 1-3 in sequential experiments.
Condition 4
$n_{tk}=n_{ck}=\frac{n_{k}}{2}$, $k=1,\ldots,K$.
Property 4
For any $k$, $\mathcal{F}_{[k]}=1-\mathcal{F}_{[k]}$.
Property 5
For any $k$ and $w\in\mathcal{F}_{[k]}$,
$P\left(W_{[k]}^{(0)}=w\right)=\frac{1}{|\mathcal{F}_{[k]}|}.$
Property 6
Based on these conditions and properties, we derive the unbiasedness and a lower bound for the variance reduction of the difference-in-means estimator under SeqVNSRR.
4.2 VNSRR for Stratified Randomized Experiments
@@
Other than rerandomization, stratification stands out as another approach to balance covariates. The idea of stratification, or blocking, was first introduced by Fisher (1926) to balance discrete covariates. In stratified randomized experiments, experimenters divide units into strata, namely, layers or levels, according to the discrete covariates and conduct complete randomization within each stratum. Wang et al. (2023) proposed two strategies that combine stratification and rerandomization to improve covariate balance further. The overall strategy rerandomizes over all strata and measures the overall covariate imbalance, while the stratum-specific strategy rerandomizes within each strata separately and measures the stratum-specific covariate imbalance. Both strategies have their unique advantages. However, since the stratum-specific strategy is identical to the setting we consider in Section 3 within each stratum, we mainly focus on the overall strategy in this section.
@@
We assume that the $n$ units are divided into $K$ strata of sizes $(n_{1},\ldots,n_{K})$. The index set of units in the $k$-th stratum is denoted by $\mathcal{I}_{k}$. In the $k$-th stratum, $n_{tk}$ units are assigned to the treatment arm and $n_{ck}=n_{k}-n_{tk}$ units are assigned to the control arm. Similar to Section 2, the assignment status is represented by a $n$-dimensional vector $W$ and the covariates are gathered into a $n\times p$ matrix $X$. The overall strategy proposed by Wang et al. (2023) samples assignments from the space $\{W\in\{0,1\}^{n}:\sum_{i\in\mathcal{I}_{k}}W_{i}=n_{tk},\forall k\}$ repeatedly until $M(W)<a$. To our knowledge, the only existing method to achieve this goal is to perform acceptance-rejection sampling, which we refer to as StratARSRR.
We note that rerandomization in stratified experiments can be formulated into the following constrained binary quadratic programming problem.
minimize $\quad M(W),$
subject to $\quad\sum_{i\in\mathcal{I}_{k}}W_{i}=n_{tk},\;k=1,\ldots,K$
$\quad W_{i}\in\{0,1\},\;i=1,\ldots,n.$ (3)
@@
Based on this observation, we extend VNSRR to stratified randomized experiments and refer to this variant as StratVNSRR. StratVNSRR performs variable neighborhood searching while considering the constraints imposed by the stratum structure. In the local search step, StratVNSRR randomly selects a list of multiple disjoint pairs of treated units and control units from each stratum. Then StratVNSRR aggregates the lists from different strata and permutes the pairs in the aggregated list. StratVNSRR follows this new ordering of unit pairs to examine whether switching each pair leads to a more balanced assignment. Compared to consecutively examining pairs in the same stratum, this procedure avoids any predetermined ordering of units or strata and allows for more flexible search in the local region. In the shaking step, StratVNSRR conducts random unit swaps within every stratum. This ensures that the new assignment is sufficiently distant from the current one.
We note that the feasible set in (4.2) is a subset of the feasible set in (3.1). As a result, StratVNSRR belongs to the class of rerandomization methods we consider in Section 3. We verify that StratVNSRR always satisfies Property 2-3 and satisfies Property 1 under Condition 4. Therefore, the statistical properties of the difference-in-means estimator under StratVNSRR can be derived as corollaries of Theorem 1 and Theorem 2.
4.3 VNSRR for Cluster Randomized Experiments
@@
Many public health and social science experiments (Donner et al., 2000; Hayes and Moulton, 2017; Schochet, 2020) assign the treatment at the cluster level rather than the individual level due to practical limitations or policy factors. Suppose the $n$ units are divided into $K$ clusters of sizes $(n_{1},\ldots,n_{K})$ in a cluster randomized experiment. We denote $\mathcal{I}_{k}$ as the index set of units in the $k$-th cluster. The experimenter assigns $K_{t}$ clusters to the treatment arm and $K_{c}=K-K_{t}$ clusters to the control arm. We note that when the $n_{i}$’s are not all equal, neither the treatment arm nor the control arm has a fixed number of units, which is different from all the settings we consider in previous sections. The assignment status of the clusters is denoted by a $K$-dimensional vector, $U$. And the assignment status of all units is denoted by a $n$-dimensional vector, $W$. $W_{i}=U_{k}$ if the $k$-th cluster includes the $i$-th unit.
@@
Lu et al. (2023) suggested conducting cluster rerandomization to balance covariates in cluster randomized experiments. The authors proposed two metrics that measure the covariate imbalance. One is the Mahalanobis distance of cluster-level covariates, and the other is the Mahalanobis distance of individual-level covariates. Since using the first metric is equal to treating a cluster as an aggregated unit and can be efficiently solved by the approach proposed in Section 3, we mainly focus on the second metric in this section. We denote the individual-level covariates as a $n\times p$ matrix, $X$. Some cluster-level covariates may also be available, but we can construct individual-level covariates from cluster-level ones by ensuring they are identical within each cluster. We then formulate cluster rerandomization into the following constrained binary quadratic programming problem.
minimize $\quad M(W),$
subject to $\quad W_{i}=U_{k},\;\text{for all }k\text{ and }i\in\mathcal{I}_{k}$
$\quad\sum_{k=1}^{K}U_{k}=K_{t},$
$\quad U_{k}\in\{0,1\},\;k=1,\ldots,K.$ (4)
We propose the ClustVNSRR method, a variant of VNSRR, to solve (4.3) efficiently and accelerate cluster rerandomization. ClustVNSRR performs variable neighborhood searching at the cluster level and treats the clusters as the operational units in the local and shaking search steps. Similar to Section 4.2, we note that the feasible set in (4.3) is a subset of the feasible set in (3.1). Therefore, we derive the statistical properties of the difference-in-means estimator under ClustVNSRR by verifying that ClustVNSRR always satisfies Property 2-3 and satisfies Property 1 under the following variant of Condition 3.
Condition 5
$K_{t}=K_{c}=\frac{K}{2}$.
@@ 5 Simulation Studies and Real Applications
This section compares VNSRR with CR, ARSRR, and PSRR. To compare the computation time of VNSRR and the existing methods, we use each method to sample 1000 acceptable assignments respectively in one simulation and replicate the simulation 100 times on a MacBook Pro (Apple M1 Max Chip, 32 GB Memory, 10 Cores). To examine the statistical properties, we use each method to sample one acceptable assignment in the design stage to conduct the experiment and 1000 acceptable assignments in the analysis stage to conduct randomization-based inference. We replicate the simulation 1000 times using a High-Performance Computing cluster.
@@ 5.1 Simple Randomized Experiments
We first consider the simple randomized experiments described in Section 3 and mimic the simulation settings in Zhu and Liu (2022). We assume that the covariates follow the standard normal distribution identically and independently, $X_{ij}\overset{i.i.d.}{\sim}N(0,1)$, $i=1,\ldots,n$, $j=1,\ldots,p$. The potential outcomes of the control group are generated independently from a linear model, $Y_{i}(0)=\sum_{j=1}^{p}X_{ij}+\epsilon_{i}$, where $\epsilon_{i}\sim N(0,p)$. We conduct simulations under both the null hypothesis and the alternative hypothesis, where we set $Y_{i}(1)=Y_{i}(0)$ and $Y_{i}(1)=Y_{i}(0)+0.2\sqrt{Var[Y_{i}(0)]}$, respectively. We consider three sample sizes, $n\in\{30,100,500\}$. The size of the treatment group and the control group are set as equal.
Morgan and Rubin (2012) proved that the asymptotic acceptance probability of the assignment is $p_{a}=P(\chi_{p}^{2}<a)$ and Li et al. (2018) recommended that $p_{a}=10^{-3}$. Wang and Li (2022) considered settings where $p_{a}$ and $p$ are allowed to grow with $n$ rather than fixed. Motivated by these works, we considered two thresholds of the Mahalanobis distance, $p_{a}\in\{10^{-3},10^{-4}\}$, and two regimes for the number of covariates, $p=2$ and $p=n/2$.
We perform Fisher randomization tests with a significance level $\alpha=0.1$ and construct randomization-based confidence intervals with a nominal coverage rate of 90%. The exact formulas of Fisher randomization test p-values and randomization-based confidence intervals are provided in the Supplement. We use the default tuning parameter of PSRR, as Zhu and Liu (2022) pointed out that different tuning parameters lead to similar performance. For each method, we examine the absolute bias and standard deviation of the difference-in-means estimator, the size and power of the Fisher randomization test, the coverage and length of confidence intervals, and the average run time to sample $10^{3}$ acceptable assignments. We also use the randomness metric in Zhu and Liu (2022) to measure the randomness of different methods in the Supplement.
Table 1 shows the statistical and computational performances of all methods in two typical settings: a low-dimensional setting with $(n,p)=(30,2)$ and a high-dimensional setting with $(n,p)=(500,250)$. The complete results are provided in the Supplement. VNSRR is significantly faster than ARSRR and PSRR in all settings. VNSRR merely requires seconds to generate 1000 acceptable assignments, even in cases where ARSRR or PSRR requires an hour. For all methods, the finite-sample biases are negligible, Fisher randomization test controls the type-1 error, and the confidence intervals are valid. Compared to CR, all the rerandomization methods reduce the variance of the difference-in-means estimator, increase the statistical power, produce tighter confidence intervals, and have comparable performances in terms of these aspects. These results demonstrate that VNSRR maintains the appealing statistical properties of rerandomization and substantially reduces the computational cost in simple randomized experiments.
Table 1: Statistical and computational performance of different methods in simple randomized experiments.
$n$ $p$ Method Statistical Inference (×$\times 10^{-3}$) Run Time (second)
Bias SD Size Power Coverage Length
30 2 CR 6.2 74 10.5 14 89.5 246 $<0.1$
@@
VNSRR 1.6 52 11.0 17 89.0 177 1.4
500 250 CR 0.8 201 10.4 73 89.6 660 $<0.1$
ARSRR 4.0 186 9.2 77 90.8 619 4137.4
PSRR 4.5 189 9.8 78 90.2 619 275.3
VNSRR 7.6 199 12.6 73 87.4 615 1.2

### v2

@@ 3.2 Statistical Properties of BRAIN
Condition 3.
@@
We prove that the difference-in-means estimator is unbiased under BRAIN.
Theorem 1.
Under Condition 3, if $W$ is generated by BRAIN, then $\mathbb{E}[\widehat{\tau}]=\tau$.
We also prove that the difference-in-means estimator has smaller variance than that under complete randomization and construct a lower bound for the variance reduction. Denote $\textnormal{V}_{\textnormal{CR}}$ as the variance of the difference-in-means estimator under complete randomization, we prove the following theorem.
Theorem 2.
Under Conditions 1-3, if $X$ has $p$ covariates and $W$ is generated by BRAIN with threshold $a$, then
$\frac{\textnormal{V}_{\textnormal{CR}}-\textnormal{Var}[\widehat{\tau}]}{\textnormal{V}_{\textnormal{CR}}}\geq\left(1-\frac{a}{p}\right)R^{2},$
where
$R^{2}=\frac{\beta^{T}\textnormal{Cov}_{\textnormal{CR}}[\overline{X}_{t}-\overline{X}_{c}]\beta}{\textnormal{V}_{\textnormal{CR}}}=\frac{n\beta^{T}S_{XX}\beta}{n_{t}n_{c}\textnormal{V}_{\textnormal{CR}}}.$
Here $\textnormal{Cov}_{\textnormal{CR}}[\overline{X}_{t}-\overline{X}_{c}]$ represents the covariance matrix of $\overline{X}_{t}-\overline{X}_{c}$ under complete randomization.
Like other rerandomization method, the difference-in-mean estimator still satisfies the same unbiasedness and variance reduction properties. This means that BRAIN is able to accelerate the process of rerandomization without sacrificing the desirable statistical properties.
4 BRAIN for Structured Randomized Experiments
In this section, we extend BRAIN to randomized experiments with various types of structures – sequential, stratified, and cluster randomized experiments. We will discuss how to adapt our algorithm to each of the three settings. The full algorithms are provided in Appendix A, as well as the validation of their unbiasedness and variance reduction properties.
4.1 BRAIN for Sequentially Randomized Experiments
In Section 3, we consider offline rerandomization, which assigns treatment to all units at once. However, in sequential experimenters where participants arrive one by one or in groups, it is common to assign treatment to units group by group in practice.
@@ 4.1 BRAIN for Sequentially Randomized Experiments
We assume that the experiment consists of $K$ stages such that the $n$ units are divided into $K$ sequential groups of sizes $(n_{1},\ldots,n_{K})$. In the $k$-th stage, $n_{t_{k}}$ units in the $k$-th group are assigned to the treatment arm, and the remaining $n_{c_{k}}$ units are assigned to the control arm. We use the subscript $[k]$ to represent the reduced version of any quantity calculated based on only the first $k$ groups. For example, $M_{[k]}(W_{[k]})$ represents the Mahalanobis distance of the covariate matrices of the first $k$ group. Sequential rerandomization aims to do the following: when the k-th group enrolls in the experiment, randomly assign treatment to the units without changing the assignment status of the previously enrolled groups, while balancing the covariates of all units in the first k groups with $M_{[k]}(W_{[k]})\leq a_{k}$.
We extend our BRAIN method to sequentially randomized experiments (denoted by SeqBRAIN, details in Algorithm 2) where it performs BRAIN in each stage of the experiment while fixing the assignments obtained in previous stages.
4.2 BRAIN for Stratified Randomized Experiments
@@
Other than rerandomization, stratification is a common alternative approach to balance covariates. In stratified randomized experiments, experimenters divide units into strata, namely, layers or levels, according to the discrete covariates and conduct complete randomization within each stratum. If stratum-specific covariate balance is of interest, we can directly apply BRAIN to each strata. If the overall covariate balance is of interest, we can modify BRAIN to effectively account for this structure.
@@
We assume that the $n$ units are divided into $K$ strata of sizes $(n_{1},\ldots,n_{K})$. The set of indices for units in the $k$-th stratum is denoted by $\mathcal{I}_{k}$. In the $k$-th stratum, $n_{t_{k}}$ units are assigned to the treatment arm and $n_{c_{k}}=n_{k}-n_{t_{k}}$ units are assigned to the control arm. Similar to Section 2, the assignment status is represented by a $n$-dimensional vector $W$ and the covariates are gathered into a $n\times p$ matrix $X$. The ’overall’ strategy proposed by [40] samples assignments from the space $\mathcal{W}_{\textnormal{strat}}=\{W\in\{0,1\}^{n}:\sum_{i\in\mathcal{I}_{k}}W_{i}=n_{t_{k}},\forall k\in[K]\}$ repeatedly until $M(W)\leq a$. To our knowledge, the only existing method to achieve this goal is to perform acceptance-rejection sampling, which we refer to as StratARSRR.
@@
We note that rerandomization in stratified experiments is equivalent to minimizing $M(W)$ under the constraint that $W\in\mathcal{W}_{\textnormal{strat}}$. We can extend BRAIN to stratified randomized experiments and refer to this variant as StratBRAIN (Algorithm 3). StratBRAIN performs local searching and perturbation iteratively while considering the constraints imposed by the stratum structure. In the local search step, StratBRAIN randomly selects a list of multiple disjoint pairs of treated units and control units from each stratum. Then StratBRAIN aggregates the lists from different strata and permutes the pairs in the aggregated list. StratBRAIN follows this new ordering of unit pairs to examine whether switching each pair leads to a more balanced assignment. Compared to consecutively examining pairs in the same stratum, this procedure avoids any predetermined ordering of units or strata and allows for more flexible search in the local region. In the perturbation step, StratBRAIN conducts random unit swaps within every stratum. This ensures that the new assignment is sufficiently distant from the current one.
4.3 BRAIN for Cluster Randomized Experiments
@@
Many public health and social science experiments [13, 18, 37] assign the treatment at the cluster level rather than the individual level due to practical limitations or policy factors. Suppose the $n$ units are divided into $K$ clusters of sizes $(n_{1},\ldots,n_{K})$ in a cluster randomized experiment. We denote $\mathcal{I}_{k}$ as the index set of units in the $k$-th cluster. Cluster randomization assigns $K_{t}$ clusters to the treatment arm and $K_{c}=K-K_{t}$ clusters to the control arm. We note that when the $n_{i}$’s are not all equal, neither the treatment arm nor the control arm has a fixed number of units, which is different from all the settings we consider in previous sections. The assignment status of the clusters is denoted by a $K$-dimensional vector, $U$. And the assignment status of all units is denoted by a $n$-dimensional vector, $W$. $W_{i}=U_{k}$ if the $k$-th cluster includes the $i$-th unit.
@@
[24] suggested conducting cluster rerandomization to balance covariates in cluster randomized experiments. The authors proposed two metrics that measure the covariate imbalance. One is the Mahalanobis distance of cluster-level covariates, and the other is the Mahalanobis distance of individual-level covariates. Since using the first metric is equal to treating a cluster as an aggregated unit and can be efficiently solved by the approach proposed in Section 3, we focus on how to modify BRAIN to minimize the second metric. We thus propose the ClustBRAIN method (Algorithm 4), a variant of BRAIN, which performs local searching and perturbation at the cluster level while treating the clusters as the operational units.
The detailed algorithms for each of the three cases are deferred to Appendix A. In all three structured settings, the difference-in-means estimator resulting from our proposed methods enjoys similar unbiasedness and variance reduction properties to those in Section 3. The details and proofs are shown in Appendix B.
@@ 5 Simulation Studies and Real Applications
This section compares BRAIN with CR, ARSRR, and PSRR. To compare the computation time of BRAIN and the existing methods, we use each method to sample 1000 acceptable assignments respectively in one simulation and replicate the simulation 100 times on a MacBook Pro (Apple M1 Max Chip, 32 GB Memory, 10 Cores). To examine the statistical properties (such as standard deviation of the treatment effect estimator and length of confidence interval), we use each method to sample one acceptable assignment in the design stage to conduct the experiment and 1000 acceptable assignments in the analysis stage to conduct randomization-based inference. We replicate the simulation 1000 times using a High-Performance Computing cluster.
We note that CR is expected to be the fastest method with the worst statistical performances as it has no requirements for covariate balance at all. ARSRR, PSRR, and BRAIN are expected to have similar performances in terms of the statistical properties because all of them belong to the rerandomization family and use the same criterion $M(W)\leq a$. Therefore, the ideal performance of BRAIN would be:
  1. 1.
    Statistical properties: similar to ARSRR and PSRR, better than CR;
  2. 2.
    Computational time: (much) faster than ARSRR and PSRR, but not CR.
@@ 5.1 Simple Randomized Experiments
We first consider the simple randomized experiments described in Section 3 and mimic the simulation settings in [46]. We assume that the covariates follow the standard normal distribution identically and independently, $X_{ij}\overset{i.i.d.}{\sim}N(0,1)$, $i=1,\ldots,n$, $j=1,\ldots,p$. The potential outcomes of the control group are generated independently from a linear model, $Y_{i}(0)=\sum_{j=1}^{p}X_{ij}+\epsilon_{i}$, where $\epsilon_{i}\sim N(0,p)$. We conduct simulations under both the null hypothesis and the alternative hypothesis:
$H_{0}:Y_{i}(1)=Y_{i}(0),$
$H_{a}:Y_{i}(1)=Y_{i}(0)+0.2\sqrt{\textnormal{Var}[Y_{i}(0)]}.$
We consider three sample sizes, $n\in\{30,100,500\}$. The size of the treatment group and the control group are set as equal.
[29] proved that the asymptotic acceptance probability of the assignment is $p_{a}=P(\chi_{p}^{2}\leq a)$ and [22] recommended setting $a$ such that $p_{a}=10^{-3}$. [41] considered settings where $p_{a}$ and $p$ are allowed to grow with $n$ rather than fixed. Motivated by these works, we consider two acceptance probabilities, $p_{a}\in\{10^{-3},10^{-4}\}$, and calculate the threshold $a$ by solving $p_{a}=P(\chi_{p}^{2}\leq a)$. We consider two regimes for the number of covariates, $p=2$ and $p=n/2$.
We perform Fisher randomization tests with a significance level $\alpha=0.1$ and construct randomization-based confidence intervals with a nominal coverage rate of 90%. For each method, we examine the absolute bias and standard deviation of the difference-in-means estimator, the size and power of the Fisher randomization test, the coverage and length of confidence intervals, and the average run time to sample 1000 acceptable assignments. We also use the randomness metric in [46] to measure the randomness of different methods. Recall that we need 1000 i.i.d. assignments to conduct randomization-based inference. If a method is almost deterministic or lacks randomness, its statistical inference performance can be impaired.
Table 1: Statistical and computational performance of different methods in simple randomized experiments. SD: standard deviation. CP: coverage probability. Len: length.
Bias SD Size Power CP Len
30 2 $10^{-3}$ CR 6.2 74 10.5 14 89.5 246 $<0.1$
@@
BRAIN 1.6 52 11.0 17 89.0 177 1.4
500 250 $10^{-4}$ CR 0.8 201 10.4 88 89.6 660 $<0.1$
PSRR 4.3 187 10.3 91 89.7 611 2878.8
BRAIN 5.7 181 9.8 92 90.2 608 1.3
Table 1 shows the statistical and computational performances of BRAIN and the other methods in two typical settings: a low-dimensional setting with $(n,p,p_{a})=(30,2,10^{-3})$ and a high-dimensional setting with $(n,p,p_{a})=(500,250,10^{-4})$. ARSRR is omitted from the high-dimensional setting since it is too slow to be a realistic option in this setting. The complete results (including ARSRR in moderately challenging settings) are provided in the Appendix D. From the results, we can see that BRAIN is significantly faster than ARSRR and PSRR in all settings. BRAIN merely requires seconds to generate 1000 acceptable assignments, even in cases where ARSRR or PSRR requires an hour. In addition, the computation time of BRAIN does not scale with the sample size $n$, covariate size $p$, or the threshold $p_{a}$ significantly as the other methods, making it an ideal algorithm to use, especially in high-dimensional settings.
For all methods, the finite-sample biases are negligible, Fisher randomization test controls the type-1 error, and the confidence intervals are valid. Compared to CR, all the rerandomization methods reduce the variance of the difference-in-means estimator, increase the statistical power, produce tighter confidence intervals, and have comparable performances in terms of these aspects. BRAIN generates assignments that are as random as those generated by ARSRR, allowing us to conduct reliable statistical inference (see Appendix D). These results demonstrate that BRAIN maintains the appealing statistical properties of rerandomization and substantially reduces the computational cost in simple randomized experiments.